Direct Materials and Direct Labor Variance Analysis

Faucet Industries Inc. manufactures faucets in a small manufacturing facility. The faucets are made from zinc. Faucet Industries has 40 employees. Each employee presently provides 36 hours of labor per week. Information about a production week is as follows:

Standard wage per hour $15.00
Standard labor time per faucet 20 min.
Standard number of lbs. of zinc 2.1 lbs.
Standard price per lb. of zinc $10.25
Actual price per lb. of zinc $10.50
Actual lbs. of zinc used during the week 19,500 lbs.
Number of faucets produced during the week 9,000
Actual wage per hour $15.50
Actual hours per week 1,440 hrs.

Required:

a. Determine the standard cost per faucet for direct materials and direct labor. Do not round intermediate calculations and round your answers to the nearest cent.

Direct materials standard cost per faucet $
Direct labor standard cost per faucet   
Total cost per faucet $

b. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative amount, and an unfavorable variance as a positive amount

Price variance $
Quantity variance   
Total direct materials cost variance $

c. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative amount, and an unfavorable variance as a positive amount.

Rate variance $
Time variance   
Total direct labor cost variance $
Answer #1
a
Direct materials standard cost per faucet 21.53 =2.1*10.25
Direct labor standard cost per faucet 5.00 =15*(20/60)
Total cost per faucet 26.53
b
Price variance 4875 Unfavorable =19500*(10.5-10.25)
Quantity variance 6150 Unfavorable =10.25*(19500-9000*2.1)
Total direct materials cost variance 11025 Unfavorable =(19500*10.5)-(9000*2.1*10.25)
c
Rate variance 720 Unfavorable =1440*(15.5-15)
Time variance -23400 Favorable =15*(1440-9000*20/60)
Total direct labor cost variance -22680 Favorable =(1440*15.5)-(9000*15*20/60)